# Exploring human-SAV interaction using large language models: The impact of psychological ownership and anthropomorphism on user experience

### v1

@@ 2.3 User experience factors
In this project, the user experience is examined from the following aspects: user’s disclosure tendency, user’s perception of the SAV’s service quality, user’s perception of enjoyment when interacting with the SAV, the sentiment of the SAV responses, and user’s acceptance of the SAVs.
2.3.1 Disclosure tendency
@@
Disclosure tendency (DT) refers to the willingness and behavior of individuals to reveal personal information to others, including computers (Zhang and Patrick Rau, 2023). During human-computer interactions, sharing personal information can bring various benefits, such as personalized services, increased convenience, enhanced safety, the establishment of intimate relationships, smoother information exchange, and ease of communication. However, it can also pose risks, including privacy concerns, identity theft, financial fraud, and unauthorized secondary use of information (Song and Kim, 2021; Cheng et al., 2022; Zhang and Patrick Rau, 2023). Interestingly, Cheng et al. (2022) found that in ride-sharing contexts, passengers tended to focus more on the benefits of disclosing personal information than on the associated risks.
@@
Moreover, Zhang and Patrick Rau (2023) observed that participants were more willing to share personal information with intelligent agents (IAs) designed to play a servant role when these agents exhibited human-like features (e.g., a friendly tone, relatable visual design) compared to IAs with robot-like characteristics (e.g., mechanical voice, robotic appearance). In the context of SAVs, users are required to disclose personal information such as their travel routes. To achieve a more personalized travel experience, users may also need to share additional preferences, such as their music tastes, to customize SAV settings. If they perceive the SAV as their own vehicle, they may be more willing to share their personal information.
Thus, it becomes essential to examine how conversational agent design influences users’ willingness to disclose personal information during interactions with SAVs, as psychological ownership and anthropomorphism prompts may shape users’ perceptions and behaviors related to disclosure tendency in distinct ways.
2.3.2 Quality of service
@@
Quality of service (QoS) is a critical factor in evaluating user experience with a product or service. Satisfied users are more likely to engage in positive word-of-mouth, exhibit a higher willingness to pay, and result in an increased intention to continue using the product or service (Dai et al., 2021; Song and Kim, 2021). Research has shown that anthropomorphism can significantly influence the quality of consumer-AI interactions, with users tending to engage for longer periods when interacting with human-like chatbots or AI devices (Wang, 2017). Additionally, quality of service has been identified as a key factor in users’ willingness to share personal information (Song and Kim, 2021) and serves as a mediator between anthropomorphic characteristics and user acceptance (Pelau et al., 2021).
2.3.3 Perceived Enjoyment
@@
Perceived Enjoyment (PE) refers to the user’s feelings of pleasure and joy during interactions with SAVs (Song and Kim, 2021). Research on human-robot interactions has shown that enjoyable interactions can lead to a greater willingness to engage and share information (Song and Kim, 2021). Similarly, in the context of human-SAV interactions, it can be assumed that if users find the SAV enjoyable to interact with, their overall user experience is likely to improve.
2.3.4 Sentiment of the SAV responses
@@
Sentiment analysis, a combination of Natural Language Processing, Computational Linguistics, and text analysis, focuses on extracting and understanding human sentiments from textual data (Bonta et al., 2019). It has been widely adopted across academia, businesses, governments, and other organizations (Wankhade et al., 2022). Sentiment analysis typically involves two classifications: subjectivity classification, which provides a subjectivity score indicating whether a text is subjective or objective, and sentiment classification, which delivers a polarity score reflecting whether a text is positive or negative (Wankhade et al., 2022; Wang et al., 2014). In the context of human-SAV interactions, these two measures (i.e., polarity and subjectivity) can be utilized to evaluate the sentiment of SAV responses. Such measures may play a key role in shaping user experiences during interactions with SAVs.
3 Method
@@
3.1 SAV User Interface design
@@
An SAV UI prototype was designed to investigate the effects of psychological ownership and anthropomorphism on user experience when interacting with SAVs. The ‘gradio’ package was used to build the UI, and the ‘openai’ package was used to access the “gpt-3.5-turbo” model using the OpenAI API in Python (Abid et al., 2019; OpenAI, 2024b). The selection of the “gpt-3.5-turbo” model for this project was guided by its conversational abilities, cost-effectiveness, and high speed, making it the ideal choice for simulating human-SAV interactions (OpenAI, 2024a). The SAV UI developed is shown in Figure 1.
Figure 1: The developed SAV user interface.
3.2 SAV prompt design

### v2

@@ II-D User experience factors
In this project, the user experience (UX) is examined from the following aspects: user’s disclosure tendency, user’s perception of the SAV’s service quality, user’s perception of enjoyment when interacting with the SAV, the sentiment of the SAV responses, and user’s acceptance of the SAVs.
II-D1 Disclosure tendency
@@
Disclosure tendency (DT) refers to the willingness and behavior of individuals to reveal personal information to others, including computers [15]. During human-computer interactions, sharing personal information can bring various benefits, such as personalized services, increased convenience, enhanced safety, the establishment of intimate relationships, smoother information exchange, and ease of communication. However, it can also pose risks, including privacy concerns, identity theft, financial fraud, and unauthorized secondary use of information [54, 55, 15]. Interestingly, [55] found that in ride-sharing contexts, passengers tended to focus more on the benefits of disclosing personal information than on the associated risks.
@@
Moreover, [15] observed that participants were more willing to share personal information with intelligent agents (IAs) designed to play a servant role when these agents exhibited human-like features (e.g., a friendly tone, relatable visual design) compared to IAs with robot-like characteristics (e.g., mechanical voice, robotic appearance). In the context of SAVs, users are required to disclose personal information such as their travel routes. To achieve a more personalized travel experience, users may also need to share additional preferences, such as their music tastes, to customize SAV settings. If they perceive the SAV as their own vehicle, they may be more willing to share their personal information. Accordingly, we examine how conversational agent design, specifically psychological ownership and anthropomorphic cues, influences users’ willingness to disclose personal information during interactions with SAVs
II-D2 Quality of service
@@
Quality of service (QoS) is a key factor in evaluating user experience with a product or service. Satisfied users are more likely to engage in positive word-of-mouth, exhibit a higher willingness to pay, and result in an increased intention to continue using the product or service [1, 54]. Research has shown that anthropomorphism can significantly influence the quality of consumer-AI interactions, with users tending to engage for longer periods when interacting with human-like chatbots or AI devices [56]. Additionally, QoS has been identified as a key factor in users’ willingness to share personal information [54] and serves as a mediator between anthropomorphic characteristics and acceptance [57].
II-D3 Perceived enjoyment
@@
Perceived Enjoyment (PE) refers to the user’s feelings of pleasure and joy during interactions with SAVs [54]. Research on human-robot interactions has shown that enjoyable interactions can lead to a greater willingness to engage and share information [54]. Similarly, in the context of human-SAV interactions, it can be assumed that if users find the SAV enjoyable to interact with, their overall user experience is likely to improve.
II-D4 Sentiment of the SAV responses
@@
Sentiment analysis, a combination of Natural Language Processing, Computational Linguistics, and text analysis, focuses on extracting and understanding human sentiments from textual data [58]. It has been widely adopted across academia, businesses, governments, and other organizations [59]. Sentiment analysis typically involves two classifications: subjectivity classification, which provides a subjectivity score indicating whether a text is subjective or objective, and sentiment classification, which delivers a polarity score reflecting whether a text is positive or negative [59, 60]. In the context of human-SAV interactions, these two measures (i.e., polarity and subjectivity) can be utilized to evaluate the sentiment of SAV responses.
III Method
@@
III-A SAV User Interface design
@@
A conversational SAV UI prototype was designed to serve as the primary interface of the conversational SAV agents for investigating the effects of psychological ownership and anthropomorphism on user experience. The ‘gradio’ package was used to build the UI, and the ‘openai’ package was used to access the “gpt-3.5-turbo” model using the OpenAI API in Python [61, 62]. The selection of the “gpt-3.5-turbo” model for this project was guided by its conversational abilities, cost-effectiveness, and high speed, making it the ideal choice for simulating human-SAV interactions [63].
III-B SAV prompt design